OpenStudy (anonymous):

the company makes a water tank in the shape of square prism. the tank is square at the base and the top. the base has side lengths 0.7m and the taank is 1.9m high. what is the voume of this tank?

OpenStudy (lgbasallote):

hint: \[A = s^2h\] where: s = length of side h = height does that help? or do you need more help?

OpenStudy (anonymous):

need more please -_-

OpenStudy (lgbasallote):

okay...what is the length of your side here?

OpenStudy (anonymous):

it says 0.7m for the side lengths of the base.

OpenStudy (lgbasallote):

right. and what is the height?

OpenStudy (anonymous):

1.9m

OpenStudy (lgbasallote):

so if you substitute those into the equation i gave \[\huge A = (0.7)^2 \times 1.9\] then simplify...what do you get?

OpenStudy (anonymous):

0.931 cubic meters. ?? so, thats it? :)

OpenStudy (lgbasallote):

i'll check first

OpenStudy (anonymous):

aww naah. that's just the area right? so how can i get the volume of it :)

OpenStudy (lgbasallote):

yup that's right

OpenStudy (lgbasallote):

that's the volume

OpenStudy (lgbasallote):

sorry...i should've written V not A

OpenStudy (anonymous):

ahh. okay ! :) thank you very much! :)

OpenStudy (lgbasallote):

that was the formula for volume

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

um. there is still a question related to that, can you help me again?

OpenStudy (anonymous):

If the tank in part a ( previous question ) is made from a single layer of sheet metal, how much sheet will be needed for the manufacture of one tank ( the tank and a top, and the base and top are also made from the same sheet metal ) -- IM REALLY CONFUSED. :( LIKE, HOW WOULD I GET THE AMOUNT OF SHEET METAL THAT IS NEEDED FOR THAT TANK. help please. :(

OpenStudy (lgbasallote):

surface area.. the surface area formula is \[\huge SA = 2s^2 + 4sh\] where: s = side length h = height got it?

OpenStudy (anonymous):

hm yup! so, is it 6.3m?

OpenStudy (anonymous):

:)

OpenStudy (lgbasallote):

no idea. i have no calculator

OpenStudy (lgbasallote):

yup. my online calculator agrees

OpenStudy (anonymous):

yey! :) hmm and the alst question related for that problem is . = the cost of the sheet metal is quite high, and the company decides that they want to limit the amount used for the tank in part A ( previous questions ) to 5.2meter squared. if the square base of the tank remains the same dimensions as before, what will be the height of the new tank? :/ thank you very much for helping me. :))

OpenStudy (lgbasallote):

so the sides will stay the same which is 0.7 \[\huge SA = 2s^2 + 4sh\] substitute the values... \[\huge 5.2 = 2(0.7)^2 + 4(0.7)h\] solve for h

OpenStudy (anonymous):

to get the "h" should i do the 2(0.7)2+4(0.7) and divide it by 5.2?

OpenStudy (anonymous):

coz im confused on how to get that h. :(

OpenStudy (anonymous):

i got 1.54m? is that right.? i divided that equation to 5.2.